# BlendScape: Enabling Unified and Personalized Video-Conferencing Environments through Generative AI

### v1

@@ 7.2. Benefits and Limitations of Composition Techniques
The participants appreciated BlendScape’s expressive power and ability to rapidly generate and iterate on environment designs with the system, creating over 300 environments across all three study tasks (115 for Task 1, 65 for Task 2, 125 for Task 3). As P13 expressed, they exerted “such minimal effort” with the system, and “[BlendScape ] does so much” to translate the prompts into a rich environment.
We also observed participants using a variety of design strategies that utilized the full range of our multimodal techniques to steer image generation. To prototype environments for the Research Paper scenario, they crafted prompts to generate settings that spark creativity (e.g., libraries, parks, coffee shops) (P1, P5-9, P11-12, P15), added objects that represent the users’ research topic (P2-3), and removed content from the environment to promote distraction-free collaboration (P10, P13, P14). To help “de-stress” the students as their deadline approached, a popular strategy was crafting meeting spaces that transport the users to relaxing or playful locations, e.g., a beach, a rainforest, and a stuffed animal-themed space (Fig. 12). Participants also used the image-to-image mode to subtly restyle the existing scene (e.g., using prompts of “warm” and “relaxing”).
Potential challenges with distracting scene elements, time & effort required to achieve the ideal design.
To enable using a system like BlendScape for professional contexts, participants expressed a need to prevent unexpected elements “that could prove to be more distracting than helpful” (P6). These ranged from minor mistakes that “look good when you first glance at it” (P5) but become more apparent upon inspecting the scene (e.g., Two Eiffel Towers in the Task 1 example environments as shown in Fig. 15) to more significant cases where the generative AI models misunderstood participants’ intent (e.g., GLIGEN inserting a playground slide rather than presentation slides for P9).
While each environment generation takes minimal effort, participants anticipated these issues requiring several iterations to fix (P1, P3-5, P7, P10, P12, P15). As P12 stated, “If we get it right, it will be good. If we get it wrong, then we’ll spend time trying to get it right.” To prevent disrupting meetings, some participants would prefer to generate personalized environments before, rather than during, a meeting (P1-3, P5).
@@ 8. Discussion
In our Discussion, we first reflect on opportunities to support distributed collaboration with BlendScape. Then, we review two challenges with BlendScape’s composition techniques and describe how to improve upon our implementation in future work. We end by discussing the limitations of our research.
8.1. Opportunities for Supporting Distributed Collaboration with BlendScape
We developed BlendScape to enable tailoring meeting environments to a wide range of distributed collaboration scenarios. All participants expressed they could envision using a system like BlendScape to generate video-conferencing environments in the future. In addition to professional contexts, for which participants preferred subtler environments that minimize distractions, they saw potential in BlendScape to support social scenarios (e.g., team-building activities, calls between friends and family members, board game nights, or baby showers).
Another goal of BlendScape was to support prior systems’ techniques to increase co-presence and improve conversation flows. It was encouraging that study participants, when using BlendScape for the first time, recognized potential collaboration affordances that blended environments could offer and prototyped examples similar to prior work (e.g., adding visuals to represent a shared task space (Hunter et al., 2014; Follmer et al., 2010), organizing users around spatial landmarks to support turn-taking in conversations (Hu et al., 2023a; Tang et al., 2023)).
While our evaluations focused on BlendScape’s expressiveness and user experience considerations, further research is needed to evaluate the impact of blended spaces on collaboration behaviors during live meetings. We outline three directions for future work, aligned with our classification of prior systems (Sec. 2.1):
(1) Capture and composition techniques for establishing a shared physical context. While prototyping the Design Brainstorming scenario, we observed that current inpainting models produced more realistic results when the capture perspective and scale of the physical task space match those of typical BlendScape environments (e.g.,, a whiteboard captured from landscape perspective). The overhead view of the physical desk often resulted in the task space being merged with the floor. Future work could study environment composition approaches for more fine-grained physical tasks that need to be recorded from a close-up view or birds-eye perspective (e.g., soldering, knitting).
(2) Abstract vs. concrete spatial metaphors for communication. BlendScape differs from other tools leveraging proxemic interactions (e.g., MirrorBlender (Grønbæk et al., 2021), OpenMic (Hu et al., 2023a)) by offering literal representations of spatial landmarks for users to organize around (e.g., chairs for sitting, walls to designate boundaries). Future comparative studies should investigate how users utilize abstract vs. detailed spaces differently and the impact on conversation flow. In particular, we anticipate challenges when users’ real-world surroundings are blended, as some of our study participants expressed discomfort at placing their videos in other users’ “personal space”.
(3) System-driven techniques for updating environments and recording collaboration history. The current BlendScape implementation saves a history of manual customizations that users make to the meeting space. As LLMs and image generation techniques advance towards operating at video rates, there could be an opportunity for a more automated approach to adapting environments to the meeting context. For example, through processing transcripts of the meeting (Chen et al., 2023), the system could infer an ideal environment design for the current meeting activity and load prior environments for discussing recurring topics.
@@ 8.2. Improving Upon BlendScape’s Composition Techniques
We developed BlendScape to achieve a balance between usability requirements for video-conferencing and technical constraints of current image generation models. Since it is infeasible to run current models at video rates, we rely on static backgrounds. However, our customization techniques (based on text-based prompts and direct manipulation) still enable users to easily and rapidly update environments to reflect the meeting progression. Our strategy of blending video streams together still allows for rendering dynamic portions of the scene by incorporating external camera feeds (as demonstrated in our Design Brainstorming scenario for conveying a physical workspace).
@@
While we expect model performance and output quality to improve in the future, we propose extensions to BlendScape to improve upon two key limitations that were surfaced in our study.
@@
(1) BlendScape can produce environments that are agnostic to the number and existing placement of users in the scene. Our current BlendScape implementation segments users from their backgrounds before performing inpainting to create blended environments. This was an intentional choice to avoid generating inappropriate or distorted depictions of humans, which is a known issue with image generation models. However, this creates a tradeoff that environments may not reflect the existing spatial layout of users in the scene (e.g., containing an overabundance of furniture for the number of users).
@@
To achieve more “people-informed” compositions, we envision using OpenPose models999ControlNet OpenPose: https://huggingface.co/lllyasviel/sd-controlnet-openpose, which perform human pose estimation, to generate furniture layouts around existing users in the scene (e.g., automatically framing them with chairs and tables). This approach would likely require multiple passes: first, an inpainting step to generate the environment, using an input image with users present to support the OpenPose model. A second inpainting pass could be performed to remove any distorted depictions of people.
@@
(2) The environment generation techniques can be overwhelmed when users’ backgrounds and text-based prompts are in contrast. Our study participants noticed that when users’ “backgrounds are so disparate, [BlendScape ] really had trouble integrating them into a believable environment” (P13). We observed this in particular with the environments in Task 2 (Fig. 16), where the four video streams had semantically different backgrounds (e.g., medieval castle vs. office space), with little relation to the prompt of Game Streaming.
@@
To achieve more cohesive blended environments, we envision “prioritizing” or preserving a larger proportion of video backgrounds that resemble the meeting prompts (e.g., by computing similarity scores between the meeting prompts and semantic descriptions of users’ video backgrounds).
8.3. Limitations
We end our Discussion by reflecting on the limitations of our work.
@@ 8.3. Limitations
Study Sample: First, we evaluated BlendScape with individuals who primarily work in user experience design and product management. Our participants’ opinions may not generalize to all potential users of BlendScape, in particular end-users outside of a design background. However, our participants had a variety of prior experience with generative AI: 7 of 15 participants use image generation models once a week or more frequently, and the remaining 8 use them once a month or not at all. This suggests that BlendScape’s composition techniques could be accessible to novice or first-time end-users.
Lack of Baseline Comparison: Given that we designed BlendScape to support a variety of meeting scenarios and a combination of distributed collaboration approaches from prior work, no singular meeting system serves as an effective baseline to compare the full range of BlendScape’s composition techniques against. Instead, we adopted a more holistic evaluation approach to demonstrate what BlendScape can enable: (1) implementing and analyzing usage scenarios, which serves as an initial comparison of BlendScape to systems from prior work; (2) studying usability considerations and customization preferences specific to BlendScape.
Ecological Validity: We pre-recorded videos with a variety of static backgrounds to simulate video calls in our study. This was to keep the environment generations relatively consistent across participants, so we could align and compare their feedback in our analysis. However, using primarily physical and dynamic video backgrounds could have elicited different design strategies and challenges from our participants (e.g., preferences towards periodically updating meeting environments to reflect changes in users’ surroundings). In future work, we plan to conduct studies using BlendScape in live meetings to investigate its affordances for supporting distributed collaborative behaviors.
Another limitation stems from using primarily virtual backgrounds in our study. Most participants chose not to use their physical environments in Task 3 (for privacy reasons or due to their surroundings being uninteresting). Although most participants did not have a first-hand experience of blending their real-world surroundings into the environment, we find it encouraging that they still brainstormed a variety of scenarios where using authentic environments could provide value, e.g., to connect family members or encourage co-workers to “share their lives in different ways” (P9).
@@
Novelty Effects: Considering the novelty of generative AI techniques, our studies around BlendScape are subject to participant response bias (Dell et al., 2012). We reduced potential bias by discussing both the benefits and limitations of BlendScape with participants and probing them with examples of poorly designed environments (e.g., some of the Game Stream environments shown in Fig. 16) to elicit potential critiques and improvements to our system.
@@ 9. Conclusion
We presented BlendScape, a system that allows meeting participants to create blended meeting spaces that are tailored to their collaboration contexts and incorporate their own physical surroundings. Through implementing scenarios and conducting evaluations with 15 end-users, we demonstrated a rich set of distributed collaboration use cases that BlendScape can enable. Participants’ feedback was encouraging in that they could rapidly express their design intentions with BlendScape and envision using similar techniques to facilitate meetings in the future. Based on their feedback, we proposed improvements to BlendScape to mitigate challenges, e.g., incohesive or distracting scene elements. Future studies around BlendScape could explore how using BlendScape during live meetings impacts collaboration behavior and whether blended environments could mitigate challenges with traditional videoconferencing tools, e.g., fatigue and lack of co-presence.
@@ Acknowledgements.
We would like to thank Amit Gulati and Henrik Turbell for their valuable feedback during the various stages of the project. We thank Sasa Junuzovic and Pat Sweeney for their valuable suggestions on working with NDI streams. Finally, we express our gratitude to study participants for their time.References
@@
- Ledo et al. (2018) David Ledo, Steven Houben, Jo Vermeulen, Nicolai Marquardt, Lora Oehlberg, and Saul Greenberg. 2018. Evaluation Strategies for HCI Toolkit Research. In Proceedings of the 2018 CHI Conference on Human Factors in Computing Systems, CHI 2018, Montreal, QC, Canada, April 21-26, 2018. ACM, 36. https://doi.org/10.1145/3173574.3173610
@@
- Sra et al. (2018) Misha Sra, Sergio Garrido-Jurado, and Pattie Maes. 2018. Oasis: Procedurally Generated Social Virtual Spaces from 3D Scanned Real Spaces. IEEE Transactions on Visualization and Computer Graphics 24, 12 (Dec. 2018), 3174–3187. https://doi.org/10.1109/TVCG.2017.2762691
@@ A.2. GPT-aided Prompt Crafting for Image Generation
For example, if the meeting activity is ”brainstorming”, and the theme of the meeting is ”mushroom forest”, then the system crafts a base prompt of Mushroom forest-themed environment for a brainstorming session. We then leverage single-shot prompting with GPT-3.5 to enhance the base prompt with contextually-relevant details:
System Prompt to GPT: Your task is to help the user create a Stable Diffusion prompt to generate an environment design. The user will specify an activity to occur in the environment and/or a theme for the space. You will provide a list of 4-5 types of objects to put in the environment and 4-5 distinct characteristics that describe the environment. The characteristics must be detailed and designed to generate visually appealing and cohesive results. Here is an example for a brainstorming activity:
Objects: "whiteboards, plants, chairs, small tables",
Environment Characteristics: "bright, open space, natural light, refreshing atmosphere, varied textures"
@@
Input Prompt (constructed by BlendScape) to GPT-3.5: Provide a list of 4-5 types of objects to put in this environment and 4-5 characteristics that describe this environment: Mushroom forest-themed environment for a brainstorming session. Return the output as comma-separated strings in JSON format: {Objects: string, Environment Characteristics: string}
The results of such a prompting strategy to GPT results in the following output.
Objects: "Mushrooms, Trees, Fairies, Moss-covered rocks, Lanterns",
Environment Characteristics": "Enchanting, Magical, Misty, Whimsical, Serene"

### v2

@@ 7.2. Benefits and Limitations of Composition Techniques
Participants appreciated the ability to rapidly generate and iterate on environment designs with BlendScape, creating over 300 scenes across all three study tasks. As P13 expressed, they exerted “such minimal effort” to specify their intentions to BlendScape, and the system “does so much” to translate their prompts into rich environments.
Participants’ design strategies utilized the full range of our multimodal techniques to steer image generation. To prototype environments for the Research Paper scenario, they added objects that represent the users’ research topic (P2-3), removed objects to promote distraction-free collaboration (P10, P13, P14), and prompted for environments they associate with creativity, e.g., parks and coffee shops (P1, P5-9, P11-12, P15). To help “de-stress” the students as their deadline approached, a popular strategy was transporting the users to relaxing or playful locations, e.g., a beach, a rainforest, and a stuffed animal-themed space (Fig. 12). Participants also used the image-to-image mode to subtly restyle the existing scene (e.g., using prompts of “warm” and “relaxing”).
Challenges with distracting scene elements, time & effort required to achieve the ideal design.
To enable using a system like BlendScape for professional contexts, participants expressed a need to prevent unexpected elements “that could prove to be more distracting than helpful” (P6). These ranged from minor mistakes that “look good when you first glance at it” (P5) but become more apparent upon closer inspection (e.g., two Eiffel Towers in Fig. 15), to more significant cases where the generative AI models misunderstood participants’ intent (e.g., GLIGEN inserting a playground slide rather than presentation slides for P9).
While each environment generation takes minimal effort, fixing distracting elements required several iterations (P1, P3-5, P7, P10, P12, P15). This impacted when and for which scenarios participants envisioned using generative AI to personalize meeting environments. For professional scenarios, some participants would prefer to customize meetings spaces before, rather than during, live calls (P1-3, P5), as the “the uncertainty of the visualizations might detract” from work discussions (P11). Scenarios with time pressure may also call for pre-meeting customizations: P12 expressed that their “therapist charges by the minute, so I don’t want to spend time doing this during a session.”
@@ 8. Discussion
The scenarios we prototyped with BlendScape (Sec. 5) and our exploratory user study (Sec. 6) demonstrated our system’s ability to enable a wide range of meeting environment designs. Overall, all participants expressed they could envision using AI image generation techniques to customize video-conferencing environments in the future, given further controls to prevent unrealistic visuals.
However, further research is required to operationalize similar customization techniques in live video-conferencing systems and understand how to effectively utilize them to achieve different collaboration goals. We discuss two avenues extending our work with BlendScape: (1) studying the new affordances for distributed collaboration that generative AI-enabled environment composition techniques could enable; (2) technical improvements to address limitations of current-day image generation models.
8.1. New Opportunities to Facilitate Collaboration through Generative AI-enabled Environment Customization
While using our classification of prior work’s environment design strategies to guide the implementation of BlendScape (Sec. 2.1), we observed new ways for these design strategies to manifest when using generative AI. For example, BlendScape creates literal representations of spatial landmarks (e.g., chairs to establish user layouts, walls to designate boundaries) as opposed to abstract proxemic metaphors in other tools (e.g., the Virtual Floor in OpenMic (Hu et al., 2023a)). Additionally, BlendScape enables ambient representations of task spaces (e.g,. rendering physical sketches in the corner of the screen in our Design Brainstorming scenario), as opposed to traditional screen-sharing capabilities which fully direct users’ attention towards shared content.
In future work, we would find it interesting to deploy BlendScape in live meeting scenarios to study the impact of these different environment designs on users’ collaboration processes, as compared to manually-crafted spaces from prior systems.
We also envision extensions to BlendScape’s implementation to enable asymmetric environments and system-driven adaptations. Per-user, rather than global, customization controls could allow users to tailor meeting spaces to their personal needs (e.g., some users may prefer distraction-free spaces, while others may focus better with visual stimuli in the environment). However, these asymmetric environments should be carefully designed to maintain consistency across users when required for the collaboration scenario (e.g., preserving user layouts for turn-taking). Incorporating real-time summarization of meeting transcripts (Chen et al., 2023) could enable making the environment a more active partner in collaboration, e.g., automatically adjusting user layouts to transition to new activities or foreshadowing upcoming topics with related visuals.
@@ 8.2. Improving Upon BlendScape’s Composition Techniques
We developed BlendScape to achieve a balance between usability requirements for video-conferencing and technical constraints of current image generation models. At the time of research, it was infeasible to run image generation models at video rates; as such, we generate environments from a single frame of users’ video backgrounds. However, our approach still enables users to rapidly update environments (via text-based prompts and direct manipulation of video feeds) and render dynamic content by blending in external camera feeds (as demonstrated in our Design Brainstorming scenario, which incorporates a live feed of a physical workspace).
@@
While we expect model performance and output quality to improve in the future, we propose extensions to BlendScape to address two key limitations that were surfaced in our study.
@@
(1) BlendScape can produce environments that ignore the number and existing placement of users in the scene. Currently, BlendScape segments and removes users from their video backgrounds before sending the backgrounds as input to Stable Diffusion (Fig. 8). This was an intentional choice to avoid generating unrealistic depictions of humans, which is a known issue with image generation models. However, this can result in environments that do not reflect the existing spatial layout of users in the scene (e.g., not containing enough furniture to seat all users).
@@
To achieve more “people-informed” compositions, we envision using OpenPose models777ControlNet OpenPose: https://huggingface.co/lllyasviel/sd-controlnet-openpose, which perform human pose estimation, to generate furniture layouts aligned with users in the scene. This approach would likely require multiple passes: first, an inpainting step to generate the environment, using input images with users present to support the OpenPose model. A second inpainting pass could detect and correct any distorted representations of people.
@@
(2) The environment generation techniques can be overwhelmed when users’ backgrounds and text-based prompts are in contrast. Our study participants noticed that when users’ “backgrounds are so disparate, BlendScape really had trouble integrating them into a believable environment” (P13). We observed this in particular with Task 2 (Fig. 16), where the four video streams had semantically different backgrounds (e.g., medieval castle vs. office space), with little relation to the prompt of Game Streaming.
@@
To achieve more cohesive blends, one solution is to mask a larger proportion of users’ video backgrounds that are relevant to the meeting prompts, thereby prioritizing these backgrounds in the resulting environment. This relevance metric could be computed by comparing the similarity between the text-based prompts and semantic descriptions of the video backgrounds.
9. Limitations
Lack of baseline comparison: Given that BlendScape introduces a new way of composing video-conferencing environments through real-time generative AI techniques, we first sought to understand the potential value that the system could provide to users through an exploratory study. The meeting scenarios we prototyped (Sec. 5) provide an initial comparison of BlendScape to existing customization suites (e.g., Together Mode, Ohyay11footnotemark: 1), as we demonstrated that BlendScape can be used to implement a majority of design strategies offered by these tools. While conducting a controlled baseline comparison was out of scope for our work, it would be a promising avenue for future research to understand requirements for manual vs. automated authoring workflows.
Generalizability of results to live meeting scenarios: Some aspects of our study design intended to ensure consistency across participants and support their agency during the study may limit the generalizability of our results to real meeting scenarios. In Tasks 1 & 2, we used pre-recorded videos with static backgrounds to generate similar environments for all participants, enabling us to compare their feedback and extract themes. However, dynamic video backgrounds could have elicited different design strategies from our participants (e.g., incorporating physical context changes to give collaborators awareness of each others’ background activities). In Task 3, most participants chose to use virtual rather than real-world backgrounds, due to their physical surroundings being uninteresting or to preserve privacy. We find it encouraging that these participants still brainstormed a variety of scenarios where incorporating authentic environments could provide value, e.g., to connect family members or encourage co-workers to “share their lives in different ways” (P9). Finally, we simulated multi-user scenarios via pre-recorded videos and prompted participants to envision how generative AI techniques could be used to support collaboration. Future studies with multiple participants should be conducted to surface design strategies and challenges specific to collaborative environment composition.
@@ 9. Limitations
Study sample and novelty effects: Our participants’ insights may not generalize to all future users of systems like BlendScape, as we primarily studied with individuals from a UX design and product management background. However, our participants’ diverse experience using image generation models suggests that BlendScape’s composition techniques are still accessible to novice users.
@@
Considering the novelty of generative AI techniques, our studies around BlendScape are subject to participant response bias (Dell et al., 2012). We reduced potential bias by discussing both the benefits and limitations of BlendScape with participants and probing them with examples of poorly designed environments (e.g., in Fig. 16) to elicit critiques and improvements to our system.
@@ 10. Conclusion
We presented BlendScape, a rendering and composition system for video-conferencing participants to tailor meeting environments to their collaboration contexts, by leveraging AI image generation techniques. Through implementing scenarios and conducting evaluations with 15 end-users, we demonstrated a rich set of meeting spaces that BlendScape can enable. Participants’ feedback was encouraging in that they could rapidly express their design intentions with BlendScape and envisioned using similar techniques to facilitate meetings in the future. Based on their feedback, we proposed improvements to BlendScape to mitigate challenges with incohesive or distracting scene elements. Future studies around BlendScape could explore the impact of its customization techniques on users’ collaborative processes and system-driven approaches to adapting environments (e.g., facilitating activity transitions with visuals representing future topics).
@@ Acknowledgements.
We thank Amit Gulati & Henrik Turbell for their valuable feedback, Sasa Junuzovic & Pat Sweeney for their advice on working with NDI streams, and Michael Nebeling & Janet Johnson for their support throughout the project. We thank our Microsoft Research colleagues and fellow interns who participated in our video and figures. Finally, we express our gratitude to our study participants for their time.References
@@
- Han et al. (2023) Dongqi Han, Denise Y. Geiskkovitch, Ye Yuan, Chelsea Mills, Ce Zhong, Amy Yo Sue Chen, Wolfgang Stuerzlinger, and Carman Neustaedter. 2023. Dr.’s Eye: The Design and Evaluation of a Video Conferencing System to Support Doctor Appointments in Home Settings. In Proceedings of the 2023 CHI Conference on Human Factors in Computing Systems, CHI 2023, Hamburg, Germany, April 23-28, 2023, Albrecht Schmidt, Kaisa Väänänen, Tesh Goyal, Per Ola Kristensson, Anicia Peters, Stefanie Mueller, Julie R. Williamson, and Max L. Wilson (Eds.). ACM, 343:1–343:18. https://doi.org/10.1145/3544548.3581350
@@
- Jr. (2007) Dan R. Olsen Jr. 2007. Evaluating user interface systems research. In Proceedings of the 20th Annual ACM Symposium on User Interface Software and Technology, Newport, Rhode Island, USA, October 7-10, 2007. ACM, 251–258. https://doi.org/10.1145/1294211.1294256
@@
- Li and Bansal (2023) Jialu Li and Mohit Bansal. 2023. PanoGen: Text-Conditioned Panoramic Environment Generation for Vision-and-Language Navigation. arxiv (2023).
@@
- Mu et al. (2024) Qianqian Mu, Marcel Borowski, Jens Emil Sloth Grønbæk, Susanne Bødker, and Eve E. Hoggan. 2024. Whispering Through Walls: Towards Inclusive Backchannel Communication in Hybrid Meetings. In Proceedings of the CHI Conference on Human Factors in Computing Systems, CHI 2024, Honolulu, HI, USA, May 11-16, 2024. ACM, 1032:1–1032:16. https://doi.org/10.1145/3613904.3642419
@@
- Numan et al. (2024) Nels Numan, Shwetha Rajaram, Balasaravanan Thoravi Kumaravel, Nicolai Marquardt, and Andrew D. Wilson. 2024. SpaceBlender: Creating Context-Rich Collaborative Spaces Through Generative 3D Scene Blending. In Proceedings of the 37th Annual ACM Symposium on User Interface Software and Technology, UIST 2024, Pittsburgh, PA, USA, October 13-16, 2024. ACM. https://doi.org/10.1145/3654777.3676361
@@ A.2. GPT-aided Prompt Crafting for Image Generation
For example, if the meeting activity is ”brainstorming”, and the theme of the meeting is ”mushroom forest”, then the system crafts a base prompt of Mushroom forest-themed environment for a brainstorming session.
We then leverage single-shot prompting with GPT-3.5 to enhance the base prompt with contextually-relevant details:
System Prompt to GPT: Your task is to help the user create a Stable Diffusion prompt to generate an environment design.
The user will specify an activity to occur in the environment and/or a theme for the space. You will provide a list of 4-5 types of objects to put in the environment and 4-5 distinct characteristics that describe the environment. The characteristics must be detailed and designed to generate visually appealing and cohesive results. Here is an example for a brainstorming activity:
Objects: ‘‘whiteboards, plants, chairs, small tables’’
Environment Characteristics:
‘‘bright, open space, natural light, refreshing atmosphere, varied textures’’
@@
Input Prompt (constructed by BlendScape) to GPT-3.5: Provide a list of 4-5 types of objects to put in this environment and 4-5 characteristics that describe this environment: Mushroom forest-themed environment for a brainstorming session. Return the output as comma-separated strings in JSON format: {Objects: string, Environment Characteristics: string}.
The results of such a prompting strategy to GPT results
in the following output.
Objects: ‘‘Mushrooms, Trees, Fairies, Moss-covered rocks, Lanterns’’
Environment Characteristics:
‘‘Enchanting, Magical, Misty, Whimsical, Serene’’